OpenStudy (anonymous):

i need help on math im a 5th grader and i was retained in 5th grade i need help for my school and for my next year to pass school if you can send some 5th grade math questions please do so thankyou and have a great day

OpenStudy (anonymous):

Type the questions out. :)

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

You want some math questions?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

\[5 - 9 = \]

OpenStudy (anonymous):

like fast:D

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

seriously i know thAT LIKE SOME HARD QUESTIONS

OpenStudy (anonymous):

5*8=? 9*9= 12* ? =24

OpenStudy (anonymous):

Do you know those?

OpenStudy (anonymous):

Well that could be a hard question for some people. IDK what you have and have not learned, therefore i do not know what you consider "hard" and 'easy",

OpenStudy (anonymous):

5X8=4O 9X9=81

OpenStudy (anonymous):

I DONT KNOW THE THIRD ONE THOUGH

OpenStudy (anonymous):

and what's 12*?=24?

OpenStudy (anonymous):

PEASANT1

OpenStudy (anonymous):

Ok now we have something to work with!

OpenStudy (anonymous):

So, It's basically algebra! you're trying to find what the question mark is...

OpenStudy (anonymous):

:d YES SIR WE DO

OpenStudy (anonymous):

What's 24/12?

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

12* ? =24 Its like we say: \[12 \times y = 24 \] SOLVE FOR Y

OpenStudy (anonymous):

288 PEASENT1

OpenStudy (anonymous):

RIGHT?

OpenStudy (anonymous):

:P I didn't want to confuse her with letters. and that's not right... the answer's 2

OpenStudy (anonymous):

24 divided by 12

OpenStudy (anonymous):

12* ? =24 Its like we say: \[12 \times y = 24 \] SOLVE FOR Y 12 * y = 24 DIVIDE BY 12 ON BOTH SIDES 12 * Y = 24 -- ---- 12 12 Y = 2

OpenStudy (anonymous):

OH OK

OpenStudy (anonymous):

So ? would equal two.

OpenStudy (anonymous):

I DIDNT UNDERSTAND IT PLEASE SEND ME MAIL ON THIS SITE BECAUSE I SEE TOO MUCH TYPING

OpenStudy (anonymous):

lets try plugging two in for the question mark.

OpenStudy (anonymous):

12*2= ?

OpenStudy (anonymous):

12 times what equals 24?

OpenStudy (anonymous):

What's 12 times 2?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Good! So we know that the question mark is two!!

OpenStudy (anonymous):

I DIDNT UNDERSTAND IT PLEASE SEND ME MAIL ON THIS SITE BECAUSE I SEE TOO MUCH TYPING

OpenStudy (anonymous):

You just did algebra! :D

OpenStudy (anonymous):

12 * 1 = 12 12 * 2 = 24 12 * 3 = 36 etc.

OpenStudy (anonymous):

I DIDNT UNDERSTAND IT PLEASE SEND ME MAIL ON THIS SITE BECAUSE I SEE TOO MUCH TYPING

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

^ stop spamming

OpenStudy (anonymous):

@some_someone Do you want to take this?

OpenStudy (anonymous):

^ @lesley12345

OpenStudy (anonymous):

YEAH?

OpenStudy (anonymous):

Why would u want mail?

OpenStudy (anonymous):

BECAUSE THERES TO MANY PEOPLE WRITING ANSWERS I CANT CONCENTRATE

OpenStudy (anonymous):

Too many people? No.

OpenStudy (anonymous):

WE "THE TOO MANY PEOPLE" ARE TRYING TO HELP YOU.

OpenStudy (anonymous):

I KNOW

OpenStudy (anonymous):

BUT THE SCREEN IS MOVING UP AND DOWN I CANT GET THE ANSWER